OpenStudy (anonymous):

find the first derivative of ((x^3)/3)-x^2

OpenStudy (anonymous):



OpenStudy (anonymous):

\[y=\frac{ x^3 }{ 3 }-x^2\] \[\frac{ d }{ dx }\left( a x^n \right)=a*nx ^{n-1}\]

OpenStudy (solomonzelman):

yes, term by term. apply the power rule:

OpenStudy (anonymous):

wouldn't you have to use the quotient rule?

OpenStudy (solomonzelman):

no

OpenStudy (solomonzelman):

you re multiplying x^2 times 1/3

OpenStudy (solomonzelman):

I mean x63 times 1/3

OpenStudy (solomonzelman):

x^3*

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle \frac{d}{dx}\left[ \frac{x^3}{3}- x^2\right]~~~\Rightarrow~~~ \frac{d}{dx}\left[ \frac{x^3}{3}\right]- \frac{d}{dx}\left[ x^2\right]\\ ~~\\~~\\ \displaystyle ~~~\Rightarrow~~~ 3\cdot\frac{d}{dx}\left[ x^3\right]~~~+~~~(-1)\cdot \frac{d}{dx}\left[ x^2\right] }\)

OpenStudy (jdoe0001):

keep in mind that \(\bf \cfrac{x^3}{3}-x^2\implies \cfrac{1}{3}\cdot x^3-x^2\) and for that, allyou need, as already pointed out, the power rule